# Mathematical Language Models: A Survey

### 14 Dec 2023 v2

@@ 5.1.3 Theorem Proving
Current research delves into exploring mathematical theorem datasets (e.g., INT [wu2020int], Feit-Thompson [huang2018gamepad], and IsarStep [li2020isarstep]) to develop novel machine learning-based theorem-proving strategies. For instance, the format of IsarStep is shown in 14. Huang et al. [huang2018gamepad] introduced the Feit-Thompson dataset, encompassing 1,602 lemmas, expanding into 83,478 proof states for the Feit-Thompson theorem [gonthier2013machine]. IsarStep [li2020isarstep] constitutes a non-synthetic dataset extracted from the largest repository of proofs handwritten by human experts in a theorem prover. Furthermore, NaturalProofs [welleck2021naturalproofs] comprises 32,000 theorem statements and proofs, 14,000 definitions, and 2,000 other types of pages (e.g., axioms, corollaries). This dataset draws from three domains: broad-coverage data sourced from ProofWiki, deep-coverage data from the Stacks project, and low-resource data extracted from mathematics textbooks.
Figure 14: An example of IsarStep, taken from [li2020isarstep].
@@
Several formal mathematical libraries [megill2019metamath, rudnicki1992overview, grabowski2015four] encompass a range of formal languages tailored for theorem proving (TP). Examples include MML [grabowski2015four], Coq [bertot2013interactive], Lean [han2021proof], Isabelle [wenzel2008isabelle], and Hol light [slind2008brief]. For instance, MML [grabowski2015four], also known as the Mizar Mathematical Library, was established to systematically construct a centralized, reusable knowledge base reflecting the standard foundations of mathematics, namely classical first-order logic and set theory. Yang et al. [yang2019learning] compiled the CoqGym dataset, comprising 71,000 human-written proofs from 123 projects developed using the Coq proof assistant. LeanStep [han2021proof] serves as the tactic proof dataset for the Lean theorem prover, featuring high-level human-written tactics alongside kernel-level proof terms. LISA [jiang2021lisa], one of the largest proof corpora for interactive theorem provers, contains 183,000 theorems and 2.16 million proof steps. This dataset is extracted from the Archive of Formal Proofs and the standard library of Isabelle. Kaliszyk et al. [kaliszyk2016holstep] introduced the HolStep dataset rooted in Higher-Order Logic (HOL) proofs. HOList [bansal2019holist] includes almost 30,000 theorems and proofs across three corpora: core, complex, and flyspeck. The core corpus contains fundamental theorems necessary for defining tactics, while the complex corpus consists of theorems related to complex calculus. Flyspeck contains the majority of the lemmas and theorems related to the Kepler conjecture [hales2017formal].
5.2 Benchmark Datasets
@@
Various benchmark datasets have been introduced to assess the performance of different mathematical algorithms, such as MAWPS [koncel2016mawps]. Huang et al. [huang2016well] developed the Dolphin18K dataset, a large-scale and diverse collection containing over 18,000 annotated math word problems. The Mathematics dataset [saxton2018analysing] encompasses problems spanning arithmetic, algebra, probability, and calculus. These problems involve sequential questions and answers presented in a free-form textual input/output format. For assessing arithmetic understanding, Mishra et al. [mishra2022numglue] introduced NumGLUE, a multi-task benchmark that evaluates AI systems across eight distinct tasks. miniF2F [zheng2021minif2f] is a dataset designed to create a unified cross-system benchmark for neural theorem proving. It comprises 488 formal Olympiad-level mathematics problem statements, encompassing Metamath, Lean, Isabelle, and HOL Light.
@@
Several datasets consider varying difficulty levels across educational stages, ranging from elementary school [miao2020diverse, shi2022language] to high school [hendrycks2021measuring]. Miao et al. [miao2020diverse] introduced the ASDiv corpus, comprising 2,305 math word problems (MWPs) categorized by problem type and elementary school grade level. The Multilingual Grade School Math Benchmark (MGSM) [shi2022language] leverages problems from GSM8K, translating them into 10 languages with the assistance of human annotators. The MATH dataset [hendrycks2021measuring] contains 12,500 problems sourced from high school math competitions. Each problem within MATH is equipped with a step-by-step solution, enabling models to learn to generate answer derivations and explanations.
@@
Furthermore, several datasets serve to evaluate foundation models [zhong2023agieval, yuan2023well]. AGIEval [zhong2023agieval] functions as a human-centric benchmark tailored to assess foundation models’ general abilities in tasks related to human cognition and problem-solving. This dataset is collected from standardized human-centric exams like college entrance exams, law school admission tests, math competitions, and lawyer qualification tests. Yuan et al. [yuan2023well] introduced the arithmetic dataset MATH 401, designed to test the latest large language models such as GPT-4, ChatGPT, InstrctGPT, Galactica, and LLaMA. It features various arithmetic expressions and offers a comprehensive analysis of the capabilities of large language models.
5.3 Augmented Datasets
@@
Augmented datasets serve to enhance existing datasets by incorporating additional samples or information. Roy et al. [roy2018mapping] introduced Aggregate by augmenting the AllArith dataset with 661 word problems from Perturb. Perturb contains problems that were manually pruned either for not yielding the desired solution $a$ or being too dissimilar from the input problem $p$. Large-scale language models like ChatGPT and GPT-4 have been utilized to expand datasets. For instance, MetaMathQA [yu2023metamath] rephrases questions from multiple perspectives without introducing extra knowledge to bootstrap mathematical questions. The Math5K dataset [li2023camel] comprises 50,000 problem-solution pairs generated using GPT-4. These datasets leverage advanced language models to enrich the existing collection of problems and solutions.
@@
Several studies have expanded existing datasets by including supplementary information, such as code programs [austin2021program, mishra2022lila] and explanations [kim2022ept]. MathQA-Python [austin2021program] is a Python adaptation of the MathQA benchmark, comprising 23,914 problems. This dataset aims to evaluate models’ capabilities in synthesizing code from intricate textual descriptions. Lila [mishra2022lila] is a comprehensive mathematical reasoning benchmark composed of 23 diverse tasks. It extends 20 existing datasets by gathering task instructions and solutions as Python programs, offering not only correct answers but also explainable solutions. PEN [kim2022ept] focuses on providing plausible and accurate explanations for solving algebraic word problems present in three benchmark datasets: ALG514, DRAW-1K, and MAWPS. Moreover, some studies have contributed statements and proofs based on existing datasets for theorem proving [jiang2022draft, welleck2022naturalprover]. For example, miniF2F+informal [jiang2022draft] is a dataset consisting of manually curated informal statements and proofs aligned with formal statements from the miniF2F dataset. NaturalProofs-Gen [welleck2022naturalprover] adapts data from NATURALPROOFS, including theorem statements, proofs, definitions, and additional pages sourced from ProofWiki. Each proof follows a multi-step structure and references a variable number of sources.
6 Challenges and Further Directions
@@
Mathematical LLMs reside in the phenomena of hallucination and faithfulness, where the generated output may lack factual accuracy or logical grounding [rawte2023survey, ji2023survey]. This phenomenon leads to the production of erroneous or misleading mathematical results, undermining the reliability of the model’s outputs. Some studies explored to address this problem by integrating extra knowledge [he_rethinking_2022], reinforcement learning from human feedback [lightman2023lets], tools [schick2023toolformer, parisi2022talm], and verify-based methods [zhou2023solving, ling2023deductive, li2023making, zhao2023verify, he_rethinking_2022, shridhar2023screws]. However, the improvement of the hallucination phenomena is limited, which influences the trustworthiness and utility of mathematical language models in practical applications and scholarly pursuits.
@@
In particular, math problems (e.g., geometry problems [gelernter1960empirical]) involve not only textual information but also various modalities such as diagrams, graphs, or mathematical symbols [lu2021iconqa]. While existing models excel in processing textual information, interpreting and reasoning across multiple modalities concurrently remains a formidable task. Few multi-modal mathematical datasets and methods are proposed for this task [zhao2022multihiertt, lu2021iconqa]. Bridging the gap between textual representations and visual/mathematical elements necessitates robust mechanisms that can effectively capture and synthesize information from disparate sources, ultimately leading to comprehensive and accurate problem-solving capabilities. In fact, the multi-modal information in mathematics is much more complex than general multi-modal tasks like vision question answering [antol2015vqa] and image captioning [hossain2019comprehensive]. Achieving proficiency in handling multi-modal mathematical problems stands as a pivotal yet intricate objective in advancing the competency of mathematical LLMs.
@@
The uncertainty of LLMs [gawlikowski2023survey, duan2023shifting] leads to the ambiguity and variability problem of mathematical problems. While these models excel in deterministic tasks, their handling of uncertainty, such as probabilistic reasoning or dealing with incomplete or vague information, poses a significant challenge. Mathematical problems often entail nuanced interpretations, fuzzy constraints, or scenarios where a single precise solution might not exist. Several studies investigated this problem via controlled generation technologies [DBLP:conf/icml/ZhouJWCS23]. However, ensuring that LLMs can navigate and appropriately account for these uncertainties while providing accurate and contextually relevant solutions remains a complex task.
@@
It is still a challenge to evaluate mathematical LMs with robust and comprehensive evaluation metrics that adequately capture the models’ performance across various mathematical tasks. Traditional evaluation metrics in natural language processing might not effectively capture the intricacies of mathematical reasoning and problem-solving. Designing evaluation benchmarks [saxton2018analysing, hendrycks2021measuring, cobbe2021training, kushman-etal-2014-learning] and metrics [wang2021gpt, chung2022scaling, thoppilan2022lamda, yuan2023well] that encompass a wide spectrum of mathematical tasks, spanning from basic arithmetic to complex theorem proving, while accounting for linguistic fluency and mathematical accuracy, remains a significant challenge. Addressing these challenges is crucial to ascertain the reliability, efficacy, and generalizability of mathematical LMs, fostering advancements in this burgeoning field.
@@
While previous models exhibit remarkable capabilities in understanding and manipulating existing mathematical concepts, their ability to autonomously devise and rigorously prove entirely new theorems presents a formidable hurdle. The development of novel mathematical theorems demands not only profound mathematical reasoning but also creative and insightful problem-solving abilities, aspects that necessitate a deeper understanding of abstract mathematical concepts beyond what the models have been trained on. Davies et al. [davies2021advancing] applied machine learning techniques to discover potential patterns and relationships among mathematical entities. Their approach involved employing attribution techniques to comprehend these connections, leveraging these insights to foster intuition and put forward conjectures. Bridging the gap between the models’ proficiency in assimilating existing mathematical knowledge and their capability to generate novel, verifiable, and impactful theorems represents a significant frontier in leveraging LLMs for the advancement of mathematical knowledge and discovery.
@@
While mathematical LMs exhibit promising potential in autonomously solving math problems, their deployment in educational settings as teaching aids or tutors necessitates addressing several pivotal challenges. Tailoring these models to serve as effective educational tools demands not only mathematical proficiency but also adeptness in pedagogy and instructional methodologies. Few studies applied Socratic questioning for mathematical teaching [zeng2022socratic, shridhar2022automatic]. Customizing LLMs to cater to diverse learning styles, adapting explanations to suit different proficiency levels, and fostering an interactive and engaging learning environment is interesting.
@@
The training data significantly influences the performance of language models, particularly in LLMs [ding_enhancing_2023]. High-quality and diverse training data can assist the model in enhancing its mathematical reasoning capabilities [luo2023wizardmath]. As discussed in Section 4.1, while there have been limited studies on constructing instruction data through LLMs, these efforts have only considered building from a small set of mathematical reasoning datasets, such as GSM8k [cobbe2021training] and MATH [hendrycks2021measuring]. High-quality and diverse mathematical instruction data remains scarce. There is a necessity to explore and investigate additional forms and construction methods for mathematical train data. Additionally, the generation of mathematical training datasets in a multimodal context is also a promising direction.
7 Conclusions

### 23 Feb 2024 v3

@@ 5.1.3. Theorem Proving
Current research delves into exploring mathematical theorem datasets (e.g., INT (Wu et al., 2021), Feit-Thompson (Huang et al., 2019), and IsarStep (Li et al., 2021)) to develop novel machine learning-based theorem-proving strategies. For instance, the format of IsarStep is shown in 14. Huang et al. (Huang et al., 2019) introduced the Feit-Thompson dataset, encompassing 1,602 lemmas, expanding into 83,478 proof states for the Feit-Thompson theorem (Gonthier et al., 2013). IsarStep (Li et al., 2021) constitutes a non-synthetic dataset extracted from the largest repository of proofs handwritten by human experts in a theorem prover. Furthermore, NaturalProofs (Welleck et al., 2021) comprises 32,000 theorem statements and proofs, 14,000 definitions, and 2,000 other types of pages (e.g., axioms, corollaries). This dataset draws from three domains: broad-coverage data sourced from ProofWiki, deep-coverage data from the Stacks project, and low-resource data extracted from mathematics textbooks.
Figure 14. An example of IsarStep, taken from (Li et al., 2021).
@@
Several formal mathematical libraries (Megill and Wheeler, 2019; Rudnicki, 1992; Grabowski et al., 2015) encompass a range of formal languages tailored for theorem proving (TP). Examples include MML (Grabowski et al., 2015), Coq (Bertot and Castéran, 2013), Lean (Han et al., 2022), Isabelle (Wenzel et al., 2008), and Hol light (Slind and Norrish, 2008). For instance, MML (Grabowski et al., 2015), also known as the Mizar Mathematical Library, was established to systematically construct a centralized, reusable knowledge base reflecting the standard foundations of mathematics, namely classical first-order logic and set theory. Yang et al. (Yang and Deng, 2019) compiled the CoqGym dataset, comprising 71,000 human-written proofs from 123 projects developed using the Coq proof assistant. LeanStep (Han et al., 2022) serves as the tactic proof dataset for the Lean theorem prover, featuring high-level human-written tactics alongside kernel-level proof terms. LISA (Jiang et al., 2021), one of the largest proof corpora for interactive theorem provers, contains 183,000 theorems and 2.16 million proof steps. This dataset is extracted from the Archive of Formal Proofs and the standard library of Isabelle. Kaliszyk et al. (Kaliszyk et al., 2017) introduced the HolStep dataset rooted in Higher-Order Logic (HOL) proofs. HOList (Bansal et al., 2019) includes almost 30,000 theorems and proofs across three corpora: core, complex, and flyspeck. The core corpus contains fundamental theorems necessary for defining tactics, while the complex corpus consists of theorems related to complex calculus. Flyspeck contains the majority of the lemmas and theorems related to the Kepler conjecture (Hales et al., 2017).
5.2. Benchmark Datasets
@@
Various benchmark datasets have been introduced to assess the performance of different mathematical algorithms, such as MAWPS (Koncel-Kedziorski et al., 2016). Huang et al. (Huang et al., 2016) developed the Dolphin18K dataset, a large-scale and diverse collection containing over 18,000 annotated math word problems. The Mathematics dataset (Saxton et al., 2019) encompasses problems spanning arithmetic, algebra, probability, and calculus. These problems involve sequential questions and answers presented in a free-form textual input/output format. For assessing arithmetic understanding, Mishra et al. (Mishra et al., 2022b) introduced NumGLUE, a multi-task benchmark that evaluates AI systems across eight distinct tasks. miniF2F (Zheng et al., 2023) is a dataset designed to create a unified cross-system benchmark for neural theorem proving. It comprises 488 formal Olympiad-level mathematics problem statements, encompassing Metamath, Lean, Isabelle, and HOL Light.
@@
Several datasets consider varying difficulty levels across educational stages, ranging from elementary school (Miao et al., 2020; Shi et al., 2022) to high school (Hendrycks et al., 2021b). Miao et al. (Miao et al., 2020) introduced the ASDiv corpus, comprising 2,305 math word problems (MWPs) categorized by problem type and elementary school grade level. The Multilingual Grade School Math Benchmark (MGSM) (Shi et al., 2022) leverages problems from GSM8K, translating them into 10 languages with the assistance of human annotators. The MATH dataset (Hendrycks et al., 2021b) contains 12,500 problems sourced from high school math competitions. Each problem within MATH is equipped with a step-by-step solution, enabling models to learn to generate answer derivations and explanations.
@@
Furthermore, several datasets serve to evaluate foundation models (Zhong et al., 2023; Yuan et al., 2023b). AGIEval (Zhong et al., 2023) functions as a human-centric benchmark tailored to assess foundation models’ general abilities in tasks related to human cognition and problem-solving. This dataset is collected from standardized human-centric exams like college entrance exams, law school admission tests, math competitions, and lawyer qualification tests. Yuan et al. (Yuan et al., 2023b) introduced the arithmetic dataset MATH 401, designed to test the latest large language models such as GPT-4, ChatGPT, InstrctGPT, Galactica, and LLaMA. It features various arithmetic expressions and offers a comprehensive analysis of the capabilities of large language models.
5.3. Augmented Datasets
@@
Augmented datasets serve to enhance existing datasets by incorporating additional samples or information. Roy et al. (Roy and Roth, 2018) introduced Aggregate by augmenting the AllArith dataset with 661 word problems from Perturb. Perturb contains problems that were manually pruned either for not yielding the desired solution $a$ or being too dissimilar from the input problem $p$. Large-scale language models like ChatGPT and GPT-4 have been utilized to expand datasets. For instance, MetaMathQA (Yu et al., 2023b) rephrases questions from multiple perspectives without introducing extra knowledge to bootstrap mathematical questions. The Math5K dataset (Li et al., 2023a) comprises 50,000 problem-solution pairs generated using GPT-4. These datasets leverage advanced language models to enrich the existing collection of problems and solutions.
@@
Several studies have expanded existing datasets by including supplementary information, such as code programs (Austin et al., 2021; Mishra et al., 2022a) and explanations (Kim et al., 2022). MathQA-Python (Austin et al., 2021) is a Python adaptation of the MathQA benchmark, comprising 23,914 problems. This dataset aims to evaluate models’ capabilities in synthesizing code from intricate textual descriptions. Lila (Mishra et al., 2022a) is a comprehensive mathematical reasoning benchmark composed of 23 diverse tasks. It extends 20 existing datasets by gathering task instructions and solutions as Python programs, offering not only correct answers but also explainable solutions. PEN (Kim et al., 2022) focuses on providing plausible and accurate explanations for solving algebraic word problems present in three benchmark datasets: ALG514, DRAW-1K, and MAWPS. Moreover, some studies have contributed statements and proofs based on existing datasets for theorem proving (Jiang et al., 2023b; Welleck et al., 2022). For example, miniF2F+informal (Jiang et al., 2023b) is a dataset consisting of manually curated informal statements and proofs aligned with formal statements from the miniF2F dataset. NaturalProofs-Gen (Welleck et al., 2022) adapts data from NATURALPROOFS, including theorem statements, proofs, definitions, and additional pages sourced from ProofWiki. Each proof follows a multi-step structure and references a variable number of sources.
6. Challenges and Further Directions
@@
Mathematical LLMs reside in the phenomena of hallucination and faithfulness, where the generated output may lack factual accuracy or logical grounding (Rawte et al., 2023; Ji et al., 2023). This phenomenon leads to the production of erroneous or misleading mathematical results, undermining the reliability of the model’s outputs. Some studies explored to address this problem by integrating extra knowledge (He et al., 2022), reinforcement learning from human feedback (Lightman et al., 2023), tools (Schick et al., 2023; Parisi et al., 2022), and verify-based methods (Zhou et al., 2023d; Ling et al., 2023; Li et al., 2023b; Zhao et al., 2023a; He et al., 2022; Shridhar et al., 2023). However, the improvement of the hallucination phenomena is limited, which influences the trustworthiness and utility of mathematical language models in practical applications and scholarly pursuits.
@@
In particular, math problems (e.g., geometry problems (Gelernter et al., 1960)) involve not only textual information but also various modalities such as diagrams, graphs, or mathematical symbols (Lu et al., 2021b). While existing models excel in processing textual information, interpreting and reasoning across multiple modalities concurrently remains a formidable task. Few multi-modal mathematical datasets and methods are proposed for this task (Zhao et al., 2022; Lu et al., 2021b). Bridging the gap between textual representations and visual/mathematical elements necessitates robust mechanisms that can effectively capture and synthesize information from disparate sources, ultimately leading to comprehensive and accurate problem-solving capabilities. In fact, the multi-modal information in mathematics is much more complex than general multi-modal tasks like vision question answering (Antol et al., 2015) and image captioning (Hossain et al., 2019). Achieving proficiency in handling multi-modal mathematical problems stands as a pivotal yet intricate objective in advancing the competency of mathematical LLMs.
@@
The uncertainty of LLMs (Gawlikowski et al., 2023; Duan et al., 2023) leads to the ambiguity and variability problem of mathematical problems. While these models excel in deterministic tasks, their handling of uncertainty, such as probabilistic reasoning or dealing with incomplete or vague information, poses a significant challenge. Mathematical problems often entail nuanced interpretations, fuzzy constraints, or scenarios where a single precise solution might not exist. Several studies investigated this problem via controlled generation technologies (Zhou et al., 2023a). However, ensuring that LLMs can navigate and appropriately account for these uncertainties while providing accurate and contextually relevant solutions remains a complex task.
@@
It is still a challenge to evaluate mathematical LMs with robust and comprehensive evaluation metrics that adequately capture the models’ performance across various mathematical tasks. Traditional evaluation metrics in natural language processing might not effectively capture the intricacies of mathematical reasoning and problem-solving. Designing evaluation benchmarks (Saxton et al., 2019; Hendrycks et al., 2021b; Cobbe et al., 2021; Kushman et al., 2014) and metrics (Wang and Komatsuzaki, 2021; Chung et al., 2022; Thoppilan et al., 2022; Yuan et al., 2023b) that encompass a wide spectrum of mathematical tasks, spanning from basic arithmetic to complex theorem proving, while accounting for linguistic fluency and mathematical accuracy, remains a significant challenge. Addressing these challenges is crucial to ascertain the reliability, efficacy, and generalizability of mathematical LMs, fostering advancements in this burgeoning field.
@@
While previous models exhibit remarkable capabilities in understanding and manipulating existing mathematical concepts, their ability to autonomously devise and rigorously prove entirely new theorems presents a formidable hurdle. The development of novel mathematical theorems demands not only profound mathematical reasoning but also creative and insightful problem-solving abilities, aspects that necessitate a deeper understanding of abstract mathematical concepts beyond what the models have been trained on. Davies et al. (Davies et al., 2021) applied machine learning techniques to discover potential patterns and relationships among mathematical entities. Recently, Bernardino et al. (Bernardino Romera-Paredes et al., 2023) proposed FunSearch to find the first discoveries made for established open problems using LLMs. Bridging the gap between the models’ proficiency in assimilating existing mathematical knowledge and their capability to generate novel, verifiable, and impactful theorems represents a significant frontier in leveraging LLMs for the advancement of mathematical knowledge and discovery.
@@
While mathematical LMs exhibit promising potential in autonomously solving math problems, their deployment in educational settings as teaching aids or tutors necessitates addressing several pivotal challenges. Tailoring these models to serve as effective educational tools demands not only mathematical proficiency but also adeptness in pedagogy and instructional methodologies. Few studies applied Socratic questioning for mathematical teaching (Zeng et al., 2022; Shridhar et al., 2022). Customizing LLMs to cater to diverse learning styles, adapting explanations to suit different proficiency levels, and fostering an interactive and engaging learning environment is interesting.
@@
The training data significantly influences the performance of language models, particularly in LLMs (Ding et al., 2023). High-quality and diverse training data can assist the model in enhancing its mathematical reasoning capabilities (Luo et al., 2023). As discussed in Section 4.1, while there have been limited studies on constructing instruction data through LLMs, these efforts have only considered building from a small set of mathematical reasoning datasets, such as GSM8k (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b). High-quality and diverse mathematical instruction data remains scarce. There is a necessity to explore and investigate additional forms and construction methods for mathematical train data. Additionally, the generation of mathematical training datasets in a multimodal context is also a promising direction.
7. Conclusions